. An insulated tank is divided into two parts by a partition. One part of the tank contains 6 kg of an ideal gas at 50°C and 800 kPa while the other part is evacuated. The partition is now removed, and the gas expands to fill the entire tank. Determine the final temperature and the pressure in the tank.
1
Expert's answer
2020-04-23T13:13:45-0400

This is free expansion process and in this process the initial temperature and the final temperature remain same. So, "T_1=T_2= 50\degree C"


"P_1V_1=P2V_2"


"800\times \frac{V}{2}= V\times P_2"


"P_2= 400 Pa"